### v3

# CHIPS-FF: Evaluating Universal Machine Learning Force Fields for Material Properties

Abstract
In this work, we introduce CHIPS-FF (Computational High-Performance Infrastructure for Predictive Simulation-based Force Fields), a universal, open-source benchmarking platform for machine learning force fields (MLFFs). This platform provides robust evaluation beyond conventional metrics such as energy, focusing on complex properties including elastic constants, phonon spectra, defect formation energies, surface energies, and interfacial and amorphous phase properties. Utilizing 16 graph-based MLFF models including ALIGNN-FF, CHGNet, MatGL, MACE, SevenNet, ORB, MatterSim and OMat24, the CHIPS-FF workflow integrates the Atomic Simulation Environment (ASE) with JARVIS-Tools to facilitate automated high-throughput simulations. Our framework is tested on a set of 104 materials, including metals, semiconductors and insulators representative of those used in semiconductor components, with each MLFF evaluated for convergence, accuracy, and computational cost. Additionally, we evaluate the force-prediction accuracy of these models for close to 2 million atomic structures. By offering a streamlined, flexible benchmarking infrastructure, CHIPS-FF aims to guide the development and deployment of MLFFs for real-world semiconductor applications, bridging the gap between quantum mechanical simulations and large-scale device modeling.
Keywords: Machine learning force field; deep learning; foundational models; density functional theory; high-throughput; materials discovery; semiconductors
Multiscale modeling in materials science 1, 2 is an essential tool to bridge the gap between atomistic properties, device performance, and manufacturing. With the passage of the U.S. Creating Helpful Incentives to Produce Semiconductors (CHIPS) and Science Act in 2022 3, there has been a renewed focus on modeling semiconductor materials, devices, and components. Various approaches to semiconductor modeling can be applied at different scales, including quantum mechanical tools such as density functional theory (DFT)4, classical atomistic simulations such as molecular dynamics (MD) 5, technology computer-aided design (TCAD) 6 and various machine learning (ML) models 7, 8, 9, 10, 11, 12, 13.
In order to better inform semiconductor device models, accurate structure-to-property relationships must be established, which often require more computationally expensive quantum mechanical approaches. Accurate structural relaxations of defect structures and semiconductor interfaces are key to uncovering the underlying physics that govern device performance. For example, defects in semiconductors can introduce energy levels within the band gap, which can impact carrier mobility and concentration and create traps, leading to non-radiative recombination of holes and electrons 14, 15, 16, 17, 18, 19. Semiconductor surfaces and interfaces are especially important to consider since they are the building blocks of devices, impacting charge carrier behavior (band alignment), electrical properties (i.e., Schottky vs. Ohmic), thermal management and device efficiency 20, 21, 22, 23, 24, 25. In addition, the combination of semiconductor materials and high-dielectric amorphous structures are essential for building Metal-Oxide-Semiconductor Field-Effect Transistors (MOSFETs) 26, 27, 28, 29, 30, 31.
Due to the costly scaling of DFT ($N^{3}$, where $N$ is the number of electrons), high-throughput calculations of larger simulation cells required to study defects, surfaces and interfaces are not feasible. Unfortunately, many of the current classical atomistic interatomic potentials do not have the capability to achieve near-DFT accuracy and capture these complex interactions. In addition, most classical potentials are limited in their transferability to various systems and experimental conditions. Early attempts to create a universal force field for the entire periodic table were limited in their accuracy 32. For this reason, many researchers have explored machine learning force fields (MLFFs) trained on highly-accurate DFT calculations as a viable route to bridge the gap between quantum mechanical accuracy and large-scale atomistic simulations.
One of the pioneering MLFF architectures was developed by Behler and Parrinello in 2007 using neural networks 33. Other MLFF methods include Gaussian processes-based Gaussian approximation potentials (GAP) 34, spectral neighbor analysis potential (SNAP) 35 and Allegro 36. Graph neural network (GNN)-based machine learning force fields (MLFFs) are now considered state-of-the-art due to their superior accuracy and transferability. With the public availability of large DFT databases (such as the Materials Project 37, JARVIS-DFT 38, 39, OQMD 40, 41, and Alexandria 42, 43, 44, 45), several researchers have trained GNN force fields on full DFT databases, encompassing the entire periodic table. These so-called universal/unified/foundational machine learning force fields (uMLFFs) have been successful at reproducing near-DFT accuracy for a wide-range of systems.
A major advantage of utilizing these pretrained uMLFF architectures is that expensive DFT dataset generation is not an initial requirement. These pretrained uMLFFs can be used in place of DFT or fine tuned on more tailored datasets or tasks. Such uMLFFs have shown remarkable potential for numerous applications including modeling of defects and surfaces 46, 47, disordered alloys 48, high-pressure superconductors 49, solid-solution energetics and ion migration barriers 46, catalysis 50, thermal conductivity 51, semiconductor interfaces 21, guiding inverse materials design 12, 52, 53 and many other applications in chemistry, physics and materials science 54, 55, 56. Important quantities such as optimized structure, elastic properties, vibrational properties (phonons), stability, thermal properties, defect formation energy and surface energy can be easily computed from uMLFFs at a significantly lower cost than DFT. Additionally, uMLFF structural relaxations of bulk semiconductors, point defects, surfaces and interfaces can be coupled with other low-cost tools (direct prediction with GNNs 57, 17, 21, 17, 58, tight binding 59, 60) to predict electronic properties. uMLFFs can also be used to perform structure searches (replacing hundreds to tens of thousands of DFT relaxation calculations), which can be followed by a single DFT calculation for electronic properties of the target structure.
One of the first graph-based uMLFF architectures introduced in 2022 was M3GNet from Chen and Ong 61. M3GNet has since transformed into MatGL (Materials Graph Library) 62, a re-implementation of M3GNet built on the Deep Graph Library (DGL) 63 and PyTorch 64. M3GNet evolved from the MEGNet (MatErials Graph Network) property prediction model introduced in Ref. 65. M3GNet was trained on $\approx$ 60,000 inorganic materials ($\approx$ 190,000 relaxation steps) in the Materials Project (MP), covering 89 elements of the periodic table 61. This dataset is commonly referred to as MPF 61. In M3GNet, the angles (representing three-body interactions) are incorporated by aggregating to bonded atoms within the graph convolution steps to update atoms, bonds and properties 61. The Atomistic LIne Graph Neural Network (ALIGNN)-FF model 66 was introduced in 2022 and was trained on the JARVIS-DFT dataset of $\approx$ 75,000 materials ($\approx$ 300,000 relaxation steps, ALIGNN-FF DB). ALIGNN-FF evolved from the ALIGNN property prediction model released in 2021 58. The ALIGNN model achieved significant improvement in property prediction tasks due to the inclusion of bond-angles in terms of line graphs, which previous GNN models lacked 58. The earlier versions of ALIGNN-FF utilized a k-nearest neighbor graph 66, while the updated version of ALIGNN-FF (2024.12.2) utilizes a radius graph, making it faster and more robust. Another uMLFF that utilizes a crystal graph (atom graph and bond graph) is the Crystal Hamiltonian Graph Neural Network (CHGNet) 67, which was proposed in 2023. This model was trained on the Materials Project Trajectory (MPtrj) dataset 67, which includes DFT calculations for $\approx$ 150,000 materials ($\approx$ 1.58 million relaxation steps). CHGNet also incorporates magnetic moments into the training process to enhance the description of chemical reactions 67.
MACE68, 69, 54, an equivariant message passing neural network potential, uses higher order messages (higher than two body) combined with atomic cluster expansion (ACE) 70, 71, which is a method for deriving an efficient body-ordered symmetric polynomial basis to represent functions of atomic neighborhoods. A pre-trained model for the entire periodic table (trained on MPTrj) for MACE was released in late 2023 54 (referred to as MACE-MP-0). In late 2024, a new pretrained version of MACE (known as MACE-MPA-0) was released, which is trained on a combination of MPTrj and Alexandria 72. SevenNet (Scalable EquiVariance Enabled Neural Network) 73, another equivariant message passing uMLFF, is based on the NequIP 74 architecture for building E(3)-equivariant force fields. SevenNet trained on the MPtrj dataset and was released in 2024. In addition to these open source models, there have been proprietary uMLFF models developed including MatterSim (Microsoft) 75 and GNoME (Google DeepMind) 76. A version of MatterSim (MatterSim-v1), which has a similar but slightly modified architecture to M3GNet and is trained on over 17 million structures sampled from Materials Project, Alexandria, and Microsoft’s own dataset (consisting of additional structures and MD trajectories at ambient and extreme conditions), was made open source to the community in December 2024 77. The Orb 78 pretrained uMLFF, released in fall 2024, was developed by the startup company Orbital Materials, where superior performance for force and energy predictions were demonstrated at significantly reduced computational cost 55, 79, 80. Orb utilizes an attention augmented Graph Network-based Simulator (GNS) 81, which is a type of Message Passing Neural Network (MPNN) 82. In contrast to equivariant models such as MACE and SevenNet, Orb does not utilize equivariant message passing. The open source Orb model (orb-v2) is trained on MPtrj and Alexandria (over 3 million materials and 32 million relaxation steps) 55. In late 2024, Meta FAIR Chemistry released the Open Materials 2024 (OMat24) 83 models and dataset, which contains over 110 million DFT calculations and pretrained uMLFF models based on the EquiformerV2 84 architecture, which is a highly accurate equivariant transformer model. OMat24 offers various sized models trained on OMat, MPTrj and Alexandria.
Recently, interest in the field of uMLFFs has grown immensely (from academia, government and industry), as seen by the several models above and various use cases of new model architectures. Given the current state of the field, benchmarking the performance of uMLLFs is imperative. Large-scale efforts to benchmark the performance of various pretrained uMLFF architectures have emerged such as Matbench Discovery 85, 80 and the JARVIS-Leaderboard 86, 87, which calculate error metrics and allow for new models to be uploaded as time progresses. There have been focused benchmarking efforts for more involved properties beyond energy. Yu et al. 88 performed a systematic assessment of uMLFFs (M3GNet, ALIGNN-FF, CHGNet and MACE), calculating properties such as optimized structure, formation energy, bulk modulus, and phonon band structure, comparing directly to DFT to calculate error metrics. Benchmarking of more advanced properties such as surface energies, defect formation energies, solid-solution energetics, and ion migration barriers have recently been performed with M3GNet, CHGNet and MACE 46, 47. In addition, a high-level assessment of uMLFF computed thermal conductivity has been conducted with M3GNet, CHGNet, MACE, SevenNet and ORB by Pot́a et al. 51 More recently, Zhu et al. utilized M3GNet, CHGNet, MACE, SevenNet, and ORB to accelerate the prediction of CALPHAD-based phase diagrams for complex alloys 89 and Lowe et al. performed in-depth benchmarking of phonons for over 10,000 materials using all of the recent state-of-the-art uMLFFs 90. These recent efforts have highlighted the successes and limitations of various uMLFF architectures, and have emphasized the need for a comprehensive publicly available benchmarking platform. Such a platform allows the community to make informed decisions on regarding uMLFFs, outweighing factors such as computational cost versus accuracy. In addition, researchers may find systematic trends in the uMLFF results which can help others to understand the applicability of each uMLFF type. Well-understood systematic trends in DFT functionals (i.e., the underbinding and lattice constant errors of GGA) have been exploited to identify new exfoliable van der Waals (vdW) materials 91. In contrast to the Jacob’s Ladder metaphor of DFT exchange-correlation functionals 92, where the “rungs” of the ladder refer to the increasing complexity (and computational cost) and accuracy of the functionals, there (as of yet) exists no systematic assessment or generalizable statements for uMLFFs. As uMLFFs begin to replace DFT calculations, such analysis is needed to help the community better develop new force fields or fine-tune existing force fields for downstream applications.
In this work, we present CHIPS-FF (Computational High-performance Infrastructure for Predictive Simulation-based Force Fields): a generalized user-friendly open-source benchmarking package to test various uMLFFs for a number of properties beyond energy. We have implemented a streamlined workflow to use uMLFFs for structural relaxation, bulk modulus, elastic properties, point defect formation energy, surface energy, interfacial properties (work of adhesion), molecular dynamics and creation of amorphous structures, all with automated error metrics from JARVIS-DFT. In contrast to large scale benchmarking efforts such as Matbench Discovery 85, 80, which requires users to submit contributions covering enormous test sets (on the order of 200,000), CHIPS-FF allows for robust benchmarking on smaller datasets and more complex properties. Although our framework is agnostic to the type of material, we chose a set of 104 materials that are most relevant for semiconductor devices as a case study. Additionally, we evaluate force-prediction accuracy of these models for close to 2 million atomic structures from Materials Project and JARVIS-DFT databases. We intend this codebase and benchmarking dataset to benefit the MLFF community and aid in the development of future models.
[…]
$E_{\text{vac}}=E_{\text{defect}}-E_{\text{bulk}}+\mu$ (1)
where $E_{\text{defect}}$ is the total energy of the defect supercell, $E_{\text{bulk}}$ is the energy of the bulk structure (no defect) and $\mu$ is the chemical potential of the missing atom. The lowest energy crystal structure of the elemental solid (from JARVIS-DFT) is used to calculate the chemical potential with each respective uMLFF (where the structure is fully re-relaxed). Surfaces were generated for miller indices of [1, 0, 0], [1, 1, 1], [1, 1, 0], [0, 1, 1], [0, 0, 1], and [0, 1, 0], skipping any polar surfaces in the surface relaxation. Each surface contained at least 4 layers and 18 Å of vacuum. The surface energy of each non-polar surface was calculated as:
$\gamma=\frac{E_{\text{surface}}-N\cdot E_{\text{bulk}}}{2A}$ (2)
where $E_{\text{surface}}$ is the energy of the surface structure, $E_{\text{bulk}}$ is the energy of the bulk structure, $N$ is the number of bulk unit cells in the surface, and $A$ is the cross sectional area of the surface. In order to study amorphous materials, we have implemented finite-temperature molecular dynamics into our workflow to perform melt/quench simulations using Berendsen NVT (Number of particles, Volume, Temperature ensemble) dynamics within ASE. For our case study of amorphous Si, we used a timestep of 1 femtosecond and ran the simulation at 3500 K for 10 ps (melt) and 300 K for 20 ps (quench). For benchmarking calculations to compare uMLFF results to, we performed similar ab initio MD (AIMD) calculations with the Vienna Ab initio Simulation Package (VASP) code, using projector augmented wave (PAW) pseudopotentials 106, 107 and the vdW-DF-optB88 functional. In order to perform calculations for material interfaces, we utilized InterMat 21, a newly developed package for the generation and calculation of interface structures (substrate plus film). The initial interface is obtained by creating a superlattice or alternating slab junction (ASJ) structure (without vacuum padding) 108, 109 and then using the Zur algorithm 110 to obtain the best candidate interface. From here, we computed the relative alignment between the film and substrate in the in-plane (xy) direction by performing a grid search with a 0.05 fractional spacing interval with each respective uMLFF model. In addition to determining the optimal in-plane orientation between the substrate and film, this gives us a qualitative estimate of how smooth the potential energy surface is at the interface for each pretrained uMLFF. From this calculation, we also determined the work of adhesion ($W_{\text{ad}}$) at the interface with:
$W_{\text{ad}}=\gamma_{\text{film}}+\gamma_{\text{substrate}}-\gamma_{\text{interface}}$ (3)
where $\gamma_{\text{film}}$ is the surface energy of the film, $\gamma_{\text{substrate}}$ is the surface energy of the substrate and $\gamma_{\text{interface}}$ is the interfacial energy calculated as:
$\gamma_{\text{interface}}=\frac{E_{\text{interface}}-E_{\text{film}}-E_{\text{substrate}}}{A}$ (4)
where $E_{\text{interface}}$ is the total energy of the interface, $E_{\text{film}}$ is the total energy of the film, $E_{\text{substrate}}$ is the total energy of the substrate and $A$ is the cross sectional area.
Figure 1: A full schematic of the CHIPS-FF workflow.
A majority of the reference data used to compare our uMLFF results to were obtained from the JARVIS-DFT database, which contains relaxed structural data (primitive lattice vectors, volume), elastic properties (bulk modulus, elastic tensor) and phonon band structure 38, 39. Within JARVIS-DFT, there exists a database of single vacancy calculations 16 and surfaces 21 that were both used for benchmarking these uMLFFs. All of the DFT data in JARVIS was computed with the vdW-DF-optB88 111 functional, which is a modified Generalized Gradient Aprroximation (GGA) functional designed to explicitly include van der Waals (vdW) interactions. This is slightly different than other materials repositories such as the Materials Project (MPtrj), which mostly contains DFT calculations performed with the Perdew Burke Ernzerhof (PBE) 112 GGA functional. Despite mostly comparing to JARVIS-DFT results (vdW-DF-optB88), the CHIPS-FF workflow has a flexible framework to use any DFT dataset as a ground truth (i.e., the Materials Project, Alexandria, a users own DFT dataset, etc.). In addition, the platform is flexible enough to compare uMLFF results directly to available experimental values.
Table 1: The percentage of unconverged structural relaxations for bulk, surface, and vacancy calculations for each uMLFF. The green highlights which models have the highest convergence rate.
uMLFF Type Bulk Surface Vacancy alignn_ff 6.7 37.5 39.6 chgnet 0.0 1.1 6.2 eqV2_153M_omat 0.0 14.8 31.2 eqV2_31M_omat 0.0 17.0 25.0 eqV2_31M_omat_mp_salex 0.0 5.7 0.0 eqV2_86M_omat 1.9 18.2 27.1 eqV2_86M_omat_mp_salex 0.0 5.7 2.1 mace 0.0 6.8 25.0 mace-alexandria 2.9 15.9 33.3 mace-mpa 0.0 6.8 33.3 matgl 0.0 4.5 14.6 matgl-direct 3.8 6.8 12.5 mattersim 0.0 4.5 16.7 orb-d3-v2 0.0 1.1 2.1 orb-v2 1.0 0.0 0.0 sevennet 0.0 2.3 41.7
Figure 2: a) Scaling analysis of various uMLFF up to 10,000 atoms (for a supercell of Cu), b) an example JARVIS-Leaderboard entry for the force MAE of the MLEARN dataset, c) an example of the interactive error metrics within the JARVIS-Leaderboard (Pearson correlation coefficient).
Table 2: The mean absolute error (MAE) and total computational time for a variety of material properties calculated with each uMLFF type. The MAE is calculated with respect to JARVIS-DFT data. Properties include (in order) lattice constants a, b, c, volume, bulk modulus, C11 and C44 components of the elastic tensor, and phonon band structure. The computational time is measured per CPU. An interactive and up-to-date version of the table is available in the JARVIS-Leaderboard.
uMLFF Type err_a err_b err_c err_vol err_kv err_c11 err_c44 err_phonon Time Å Å Å Å3 GPa GPa GPa cm-1 seconds alignn_ff 0.087 0.102 0.137 9.95 122 196 74 150 8.8 x 104 chgnet 0.046 0.049 0.109 3.58 85 59 45 68 3.5 x 104 eqV2_153M_omat 0.036 0.040 0.111 2.90 117 70 32 49 1.6 x 102 eqV2_31M_omat 0.029 0.032 0.099 3.05 115 78 37 48 3.4 x 105 eqV2_31M_omat_mp_salex 0.028 0.030 0.096 3.17 112 47 41 48 2.7 x 105 eqV2_86M_omat 0.033 0.040 0.094 3.41 116 85 35 49 9.7 x 105 eqV2_86M_omat_mp_salex 0.027 0.030 0.100 3.14 114 45 35 48 7.2 x 105 mace 0.035 0.038 0.084 3.00 94 43 38 58 7.2 x 104 mace-alexandria 0.081 0.086 0.206 6.05 76 96 45 81 2.5 x 105 mace-mpa 0.044 0.046 0.098 3.57 110 35 34 49 8.4 x 104 matgl 0.052 0.057 0.128 3.50 83 64 41 78 1.7 x 104 matgl-direct 0.044 0.046 0.106 2.88 84 73 50 85 2.5 x 104 mattersim 0.031 0.033 0.110 3.18 110 32 32 48 5.9 x 104 orb-d3-v2 0.031 0.030 0.048 1.90 130 74 74 52 2.7 x 104 orb-v2 0.023 0.025 0.159 3.38 119 80 74 49 2.4 x 104 sevennet 0.035 0.038 0.100 3.43 103 46 37 53 5.7 x 104
Figure 3: Parity plots for a) surface energy (in J/m2) and b) vacancy formation energy (in eV) for each uMLFF type with respect to JARVIS-DFT data (MAE is given in the inset). c) Depicts the parity plot for work of adhesion (Wad in J/m2) for several material interfaces and MAE with respect to previous calculated and experimental data (taken from Ref. 113). To the right of the parity plot, we focus on the best performing Wad prediction for the Al(111)Al2O3(001) interface, showing the xy grid search and optimal in-plane interface configurations with each uMLFF.
Figure 4: The calculated radial distribution function ($g(r)$) as a function of pair separation distance computed with several uMLFF models for amorphous Si (by performing melt/quench simulations). Ab initio (AIMD) results are given as a ground truth benchmark (black dotted line) and the MAE and R2 with respect to AIMD is reported in the inset. The atomic structure for amorphous Si is given for AIMD and the best performing uMLFF models (MatterSim, eqV2_31M_omat and ORB).
Fig. 1 depicts a full schematic of the CHIPS-FF workflow. In this workflow, initial structures are taken from the JARVIS-DFT database. JARVIS-Tools is used to pull structures from the JARVIS-DFT database and generate supercell surface and defect structures prior to atomistic simulations. Various uMLFF calculators are accessed through ASE. So far, we have implemented MatGL, CHGNet, ALIGNN-FF, MACE, SevenNet, ORB, MatterSim, and eqV2 (OMat24). As new pretrained force fields are developed over time, they will be added to the workflow. In addition, CHIPS-FF offers the flexibility for the user to add their own force field or property to the pipeline. Once each uMLFF is loaded and run through the ASE calculator, a variety of properties can be computed including: optimized structure, bulk modulus (equation of state fitting), elastic tensor (elastic package), phonons (phonopy), thermal expansion (phonopy), thermal conductivity (phono3py), molecular dynamics, vacancy formation energy, surface energy, and interface properties (InterMat). As these calculations are carried out, detailed logging information is saved, including the computational timing for each stage of the calculation and whether or not each stage converged. Once these material properties are computed with each uMLFF, they are cross checked with entries in the JARVIS-DFT database and the mean absolute error (MAE) for each property is computed. For spectral quantities such as the phonon band structure, we computed the MAE for each band along the q-point path using:
$\text{MAE}=\frac{1}{N_{q}}\sum_{q\nu}\left|\omega_{q\nu}^{\text{uMLFF}}-\omega_{q\nu}^{\text{DFT}}\right|$ (5)
where $\omega_{q\nu}$ is the phonon energy (in cm-1) at each q-point and each branch index ($\nu$) (calculated with uMLFF and DFT from JARVIS-DFT) and $N_{q}$ represents the number of wavevectors along the q-path. In addition to automatically calculating the error metrics, our workflow will create entries for the JARVIS-Leaderboard in the appropriate format (.json.zip and .csv.zip files with the proper naming convention). From here, these entries can be directly uploaded to the leaderboard and compared to other benchmarks. Although we use JARVIS-DFT for the initial structures and calculation of the error metrics, our CHIPS-FF platform is flexible to use other datasets for initial structures and ground truth comparison. Our CHIPS-FF package offers easy-to-use command line tools, which can easily be parallelized to simultaneously run hundreds or even thousands of materials/uMLFFs on CPU and GPU.
For our benchmarking study, we chose a set of 104 materials most commonly found in semiconductor devices and interfaces. This test set contains metals, semiconductors and insulators to be representative of the various parts and interfaces of integrated circuits (semiconductor/insulator interfaces, semiconductor/metal contacts, etc.). Fig. S1 depicts various distributions of properties of the test set, including space group, band gap (calculated with vdW-DF-optB88), atomic number, chemical formula type, crystal system, and Wyckoff site, and dimensionality, all obtained from JARVIS-DFT. In addition, Table S1 and S2 give detailed information of the chemical composition, crystal structure and band gap for each material, where we see that the set contains 28 metals (zero band gap), 14 insulators (band gap above 3 eV) and 62 semiconductors (band gap between 0 eV and 3 eV). In addition, we see from Fig. S1 that our test set represents a diverse set of materials ideal for testing our uMLFF workflow.
When considering the use of a pretrained uMLFF, the most important considerations should be accuracy, rate of convergence, scalability, and computational cost. After running the full CHIPS-FF workflow for the test set of 104 materials, we performed an analysis of which structural relaxation calculations reached convergence (force-maximum threshold of 0.05 eV/Å within 200 steps) for bulk materials, surfaces and vacancies. This analysis is presented in Table 1, where the percentage of unconverged results is depicted for each calculation type. Aside from ALIGNN-FF, all other uMLFF have a near-perfect convergence rate for bulk structures. With regards to surface and defect calculations, the convergence rate is more variable. Interestingly, the OMat24 models significantly vary in terms of convergence for surfaces and defects depending on training data. The OMat24 models trained on OMat+MPTrj+sAlexandria have a significantly higher rate of convergence when compared to the models solely trained on OMat. ORB and CHGNet have a successful rate of convergence compared to other models, which is remarkable due to the fact that the computational cost is so low (see Table 1). This is especially important since the computational cost of optimizing surfaces and defect structures greatly outweighs bulk unit cells.
For each of the 104 structures, we performed structural relaxations, fitting of the energy vs. volume curve (equation of state), calculation of the elastic tensor, calculation of the phonon band structure, relaxation of all nonpolar surfaces (among [1, 0, 0], [1, 1, 1], [1, 1, 0], [0, 1, 1], [0, 0, 1], and [0, 1, 0]) , and relaxation of each type of point defect in the material. The computational timing and errors in the lattice constants, elastic properties and phonons for each uMLFF model are depicted in Table 2. Out of the 104 materials in our test set, there were 85 matching entries for DFT phonon calculations in JARVIS, from which the composite MAE of the phonon band structure was computed in Table 2. For most structural and phonon properties, the OMat24 models perform exceedingly well, but are the most computationally intensive. Fitting the equation of state to obtain bulk modulus and calculating the elastic tensor prove to be a more challenging task for these uMLFF models, emphasizing difficulties in modeling the potential energy landscape far from equilibrium. We found that models such as MACE-Alexandria and MatGL have superior performance in predicting the energy vs. volume curves compared to the OMat and ORB models. For contributions to the elastic tensor (C11 and C44), we also observe relatively high errors. We see that only a few models can simultaneously predict C11 and C44 with reasonable accuracy (MAE less than 50 GPa). Interestingly, we find that eqV2_31M_omat, eqV2_86M_omat and eqV2_153M_omat can predict C44 with reasonable accuracy, but have much higher prediction errors for C11, while eqV2_31M_omat_mp_salex and eqV2_86M_omat_mp_salex do a better job at predicting C11 and C44 simultaneously. We observe that MACE-MPA-0 and MatterSim give the best results for simultaneous C11 and C44 predictions.
Excellent results for structural parameters and phonons are reported for orb-v2 and MatterSim at $\approx$ 10 - 100 times less the computational cost of the OMat24 models. ORB and MatterSim offer similar performance in terms of predicting phonon band structure when compared to several equivariant uMLFFs such as MACE, OMat24 and SevenNet. This is especially interesting since ORB and MatterSim-v1 are not equivariant uMLFFs. In addition, we see improved phonon results for the newer MACE-MPA-0 when compared to MACE-MP-0. Recently, it was reported that OMat and ORB models had substantial errors in phonon predictions when compared to models such as MatterSim, MACE, CHGNet, and SevenNet for 10,000 materials 90. It is entirely possible that ORB and OMat models can predict accurate phonons for our limited test set (104 materials) but fail for a broader and more diverse set of materials.
[…]
In addition to summing the computational timing for the entire workflow, we also examined the scaling behavior of each uMLFF (on CPU). These results for supercells of Cu (up to 10,000 atoms) are depicted in Fig. 2a). Unsurprisingly, we see that OMat model scales much more drastically with system size as compared to the other uMLFF models. In addition, we find that the invariant models have more favorable scaling than the equivariant models such as MACE and SevenNet, with SevenNet having slightly better scaling than MACE. The scalability of ORB and MatterSim is a huge advantage due to their accuracy of various material property predictions. Similar scalability studies were conducted in Ref. 55 The quantities of surface energy and vacancy formation energy are difficult to predict with machine learning methods and provide a rigorous benchmarking test for uMLFF architectures. Out of the 104 materials, there exists 85 entries for surface DFT calculations and 48 entries for defect DFT calculations within JARVIS-DFT. We used these calculations to compare our uMLFF results. Fig. 3a) and Fig. 3b) depict the parity plots and corresponding MAE for surface energy and vacancy formation energy of our test sets. The very low error for surface energy (0.18 J/m2) and vacancy formation energy (0.94 eV) highlight some of the successes of recent state-of-the-art models such as OMat24, ORB, MACE-MPA-0 and MatterSim. We also observe generally good results for MACE-MP-0 and SevenNet when predicting vacancy formation energy and surface energies. The low cost of ORB, high rate of convergence and high accuracy make it a viable tool to relax larger surfaces and defect supercells.
In addition to benchmarking the properties of bulk materials, we utilized the CHIPS-FF workflow for interface calculations. Fig. 3c depicts the parity plot for work of adhesion (Wad) for various material interfaces. The MAE for Wad is computed with respect to “ground truth” data taken from Ref. 113 (from experiment and theory). As seen in Fig. 3c, predictions for Wad are extremely poor for each uMLFF type. This is not surprising due to the fact that none of these uMLFF are trained on interface data. We went on to analyze one of the best performing predictions for an Al(111)Al2O3(001) interface (see red box on Fig. 3c). From here, we performed an in-plane (xy) grid search to find the optimal energy configuration of the interface. In addition, this in-plane scan can be a test of how smooth the potential energy surface is. On the right panel of Fig. 3c we see the results of this interface scan for a number of uMLFFs. We observe a smoother potential energy surface for MACE, SevenNet, MatterSim and OMat24 models. Although the MAE is extremely high for all uMLFF models for Wad, it is the lowest for ALIGNN-FF. This can be due in part to ALIGNN-FF being trained on vdW-corrected DFT calculations and having a substantial amount of exfoliable materials in the dataset.
We chose to benchmark the performance of uMLFF models for the generation of amorphous Si (a-Si) by performing melt/quench simulations. Fig. 4 depicts a summary of these results, where the radial distribution function (RDF) is shown as a function of pair separation distance. We benchmarked these uMLFF RDF curves against the RDF curve obtained from computationally expensive AIMD simulations for a-Si. The inset of Fig. 4 depicts the MAE of each RDF curve with respect to AIMD. We observe that architectures such as MACE, SevenNet, and OMat24 (eqV2), CHGNet, ALIGNN-FF, MatterSim and ORB have excellent accuracy, while MatGL has a slightly higher MAE and a lower R2 value. The final amorphous structures for AIMD, MatterSim, eqV2 and ORB (best performing uMLFF models) are given in the inset of Fig. 4. The exceptional performance of ORB and MatterSim (both invariant uMLFFs) for a-Si is significant due to the fact that it is much less costly than other similarly performing equivariant uMLFF models. The relatively high accuracy of ALIGNN-FF with regards to a-Si is surprising due to the fact that ALIGNN-FF had inferior performance for a number of properties and is also an invariant model. A more thorough assessment of how ALIGNN-FF and other uMLFFs can model amorphous materials will be the subject of future work, which will also utilize the CHIPS-FF infrastructure. These results bring into question whether or not equivariance is necessary to achieve accurate results for amorphous materials with uMLFFs, and more rigorous benchmarking will need to be conducted to bring clarity to this.
In addition to testing these uMLFF models on a smaller and more focused dataset, we also tested on a larger and more diverse dataset for force predictions. Unlike total energy, forces are more independent of exchange-correlation functional and the accuracy of force predictions on various datasets can give valuable insight into how a particular model performs. We chose to benchmark each uMLFF on the MLEARN dataset from Ref. 114, which consists of face-centered cubic (Cu, Ni) and body-centered cubic (Li, Mo) metals and diamond group IV semiconductors (Si, Ge) which span a wide range of crystal structures and bonding environments ($\approx$ 200 - 300 data points for each element). These results are depicted in Table 3. In addition, we benchmarked the accuracy of force predictions on very large datasets that were used to train uMLFFs. These datasets included ALIGNN_FF_DB (307,000 used to train ALIGNN-FF), MPF (188,000 used to train M3GNet), and MPTrj (1.58 million used to train CHGNet, MACE, SevenNet and used in the training of ORB and OMat models). These results are depicted in Table 4. Unsurprisingly, superior accuracy is obtained across all datasets when the OMat and ORB models are used to predict forces. We also observe highly accurate forces for MatterSim and a substantial reduction in force error when going from MACE-MP-0 to MACE-MPA-0. The CHIPS-FF package has built in functions to perform these benchmarking calculations on each respective dataset, with a flexible framework to add additional datasets (such as OMat and Alexandria) and new uMLFFs as they develop over time.
Table 3: A comparison of force errors (eV/Å) and timings (s) for the mlearn DFT dataset for Mo, Li, Si, Ni, Cu, and Ge. The green color highlights the best performing model for each element.
Mo Li Si Ni Cu Ge uMLFF Type Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) alignn_ff 0.869 1.2 x 102 0.122 9.5 x 101 0.324 9.5 x 101 0.303 2.1 x 102 0.216 1.9 x 102 0.395 4.5 x 101 chgnet 0.361 9.9 x 101 0.045 1.4 x 102 0.188 3.6 x 101 0.059 3.1 x 102 0.058 2.9 x 102 0.207 3.3 x 101 eqV2_153M_omat 0.096 3.0 x 103 0.004 3.5 x 103 0.012 6.6 x 103 0.009 1.8 x 104 0.006 9.2 x 103 0.031 4.6 x 103 eqV2_31M_omat 0.098 6.9 x 102 0.004 7.9 x 102 0.016 3.2 x 103 0.010 3.4 x 103 0.005 2.1 x 103 0.033 1.1 x 103 eqV2_31M_omat_mp_salex 0.056 6.9 x 102 0.007 7.9 x 102 0.021 2.2 x 103 0.011 2.1 x 103 0.006 3.9 x 103 0.028 1.1 x 103 eqV2_86M_omat 0.094 1.8 x 103 0.004 2.0 x 103 0.014 9.3 x 103 0.009 9.9 x 102 0.006 5.5 x 103 0.032 2.7 x 103 eqV2_86M_omat_mp_salex 0.048 1.8 x 103 0.005 2.0 x 103 0.016 4.9 x 103 0.009 5.5 x 103 0.005 9.7 x 103 0.030 2.7 x 103 mace 0.292 2.6 x 102 0.035 2.8 x 102 0.149 1.9 x 102 0.055 1.8 x 103 0.059 5.5 x 102 0.203 1.9 x 102 mace-alexandria 0.343 4.3 x 102 0.025 4.7 x 102 0.202 1.1 x 103 0.099 2.2 x 103 0.055 2.1 x 103 0.161 9.4 x 102 mace-mpa 0.130 2.1 x 102 0.014 2.5 x 102 0.083 3.5 x 102 0.026 7.6 x 102 0.013 7.4 x 102 0.063 3.5 x 102 matgl 0.333 1.0 x 101 0.042 2.1 x 101 0.226 4.1 x 101 0.069 7.4 x 101 0.124 6.5 x 101 0.299 2.2 x 101 matgl-direct 0.300 3.8 x 101 0.031 3.9 x 101 0.214 4.7 x 101 0.114 1.6 x 102 0.054 1.4 x 102 0.260 4.8 x 101 mattersim 0.121 1.2 x 102 0.024 2.2 x 101 0.067 1.5 x 102 0.021 5.3 x 102 0.012 4.6 x 102 0.068 1.3 x 102 orb-v2 0.130 1.1 x 102 0.018 1.2 x 102 0.068 1.4 x 102 0.026 2.8 x 102 0.015 2.8 x 102 0.060 1.4 x 102 sevennet 0.365 1.9 x 102 0.051 1.4 x 102 0.129 1.4 x 102 0.089 8.4 x 102 0.055 3.7 x 102 0.131 1.4 x 102
Table 4: A comparison of force errors (eV/Å) and timings (s) for different datasets: ALIGNN_FF_DB, MPF, and MPTrj. The green color highlights the best performing model.
ALIGNN_FF_DB (307k) MPF (188k) MPTrj (1.58M) MLFF Type Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) alignn_ff 0.472 2.9 x 104 0.639 4.1 x 104 0.581 3.5 x 105 chgnet 0.088 3.2 x 104 0.101 1.1 x 105 0.060 7.6 x 105 eqV2_153M_omat 0.050 1.5 x 106 - - - - eqV2_31M_omat 0.051 3.0 x 105 0.046 3.5 x 105 - - eqV2_31M_omat_mp_salex 0.055 3.9 x 105 0.032 4.3 x 105 - - eqV2_86M_omat 0.052 9.2 x 105 0.043 9.3 x 105 - - eqV2_86M_omat_mp_salex 0.056 9.2 x 105 - - - - mace 0.104 5.0 x 104 0.125 1.8 x 105 0.059 1.1 x 106 mace-alexandria 0.221 1.3 x 105 0.395 2.7 x 105 - - mace-mpa 0.072 7.0 x 105 0.061 1.3 x 105 - - matgl 0.091 1.1 x 104 0.066 3.4 x 104 0.100 1.4 x 105 matgl-direct 0.104 1.7 x 104 0.095 4.2 x 104 0.118 2.7 x 105 mattersim 0.071 8.7 x 104 0.047 7.1 x 104 0.059 1.7 x 106 orb-v2 0.062 5.0 x 104 0.049 8.4 x 104 0.028 7.5 x 105 sevennet 0.092 4.4 x 104 0.081 6.0 x 104 0.041 5.0 x 105
We have introduced CHIPS-FF, an open-source benchmarking platform for MLFF architectures, which has the capability to test properties beyond the standard total energy, including forces, phonons, elastic properties, surface energy, vacancy formation energy and properties of interfaces. We benchmarked several recent state-of-the-art uMLFF architectures on a subset of 104 materials most relevant for the semiconductor industry, taking into account accuracy with respect to DFT, convergence and computational cost. As MLFFs continue to develop over time, we expect the CHIPS-FF benchmarking platform to be critical in terms of testing the quality of uMLFFs.
1 Data Availability Statement
[…]
2 Conflicts of Interest
The authors declare no competing interests.
3 Acknowledgments
All authors thank the National Institute of Standards and Technology for funding, computational, and data-management resources. This work was performed with funding from the CHIPS Metrology Program, part of CHIPS for America, National Institute of Standards and Technology, U.S. Department of Commerce. Certain commercial equipment, instruments, software, or materials are identified in this paper in order to specify the experimental procedure adequately. Such identifications are not intended to imply recommendation or endorsement by NIST, nor it is intended to imply that the materials or equipment identified are necessarily the best available for the purpose. The authors would like to acknowledge John Bonini for a fruitful discssion regarding the manuscript and software package.